OpenStudy (anonymous):

MEDAL & FAN!! Simplify the following expression, explain your steps. y = (x^2+3x-4)/(x^2+5x+4)

OpenStudy (anonymous):

@Michele_Laino @sleepyjess @KendrickLamar2014 @mayankdevnani @imqwerty @Jaynator495 @zepdrix @freckles @uri @agent0smith @inkyvoyd @hhelpplzzzz @Koikkara @xapproachesinfinity @Awolflover1 @math&ing001 @haleyelizabeth2017

OpenStudy (kendricklamar2014):

@Wendy.Ivette11714

OpenStudy (haleyelizabeth2017):

y = (x^2+3x-4)/(x^2+5x+4) We need to factor these first... \(x^2+3x-4=(x+4)(x-1)\) \(x^2+5x+4=(x+4)(x+1)\)\[y=\frac{(x+4)(x-1)}{(x+4)(x+1)}\]There is one factor on top and bottom that can cancel out, do you know what one?

OpenStudy (anonymous):

Im not sure, would it be (x+4)?

OpenStudy (anonymous):

since its the same on the top n bottom

OpenStudy (haleyelizabeth2017):

Yeah, so we are left with what?

OpenStudy (anonymous):

y = (x-1) / (x+1)

OpenStudy (haleyelizabeth2017):

Yep! You got it!

OpenStudy (anonymous):

Thank you sm!

OpenStudy (haleyelizabeth2017):

LOL you're welcome